After much difficulty you determine that the solutions to: x^2-x-4=0 are the numbers a and b. The sum of a+b equals?
Solve for the sum of a+b
Thank you for your help, it is very much appreciated.

Answer by Alan3354(69443)

After much difficulty you determine that the solutions to: x^2-x-4=0 are the numbers a and b. The sum of a+b equals?
Solve for the sum of a+b
---------------------
For ax^2 + bx + c = 0, it's -b/a
For x^2-x-4=0 it's 1/1 = 1